# SAM-R1: Leveraging SAM for Reward Feedback in Multimodal Segmentation via Reinforcement Learning

### v1

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [15, 18, 32, 20, 5, 38, 45] have achieved remarkable progress in the field of visual understanding [23, 33, 31, 19, 35], with their capabilities extending to more complex and fine-grained perception tasks like multimodal segmentation [37]. Compared to conventional segmentation methods that rely on simple categorical labels, the reasoning segmentation task [2, 14, 27] has garnered significant attention for its flexibility and practical applicability, but it also introduces substantially greater challenges. Specifically, it requires models not only to comprehend the intent behind user-provided textual queries accurately but also to perform strong logical reasoning to generate high-quality, pixel-level segmentation outputs [36, 10].
@@
LISA [14] was the first to introduce the integration of MLLMs with segmentation models via specialized tokens, demonstrating the feasibility of applying MLLMs to reasoning segmentation tasks. Building on this foundation, subsequent studies [2, 27] have adopted similar strategies, leveraging task-specific tokens generated by MLLMs to improve pixel-level segmentation performance. While these approaches are promising, they often rely heavily on large-scale annotated datasets to jointly fine-tune the language model and the segmentation decoder. This not only increases training costs but also raises the risk of catastrophic forgetting, where models perform well on in-domain data but fail to generalize to out-of-domain scenarios [7]. Furthermore, the reasoning segmentation tasks frequently involve ambiguous and complex text queries from users, which demand strong reasoning capabilities from MLLMs to accurately interpret intent and precisely localize the target segmentation regions.
Recent research has shown that reinforcement learning (RL) can significantly enhance the reasoning capabilities of large language models (LLMs) through reward-based feedback mechanisms [11]. DeepSeek-R1 [9] leverages rule-based rewards to further improve the model’s capacity for complex reasoning. This method requires the model to undergo an extensive reasoning process before producing a final answer, with rewards assigned solely based on the correctness of the final response and its adherence to a predefined output format. Such rule-based reward designs align naturally with visual understanding tasks, which often come with accurate ground-truth (GT) annotations. Inspired by this, numerous efforts [8, 40, 44, 34] have applied Group Relative Policy Optimization (GRPO) [29] to vision-language models, incorporating task-specific reward signals. For example, VLM-R1 [30] introduces both format and accuracy rewards for general vision-language tasks, and further incorporates customized rewards tailored to specific applications to mitigate reward hacking. Seg-Zero [21] expands this paradigm by designing a more comprehensive reward system, including reasoning-format, segmentation-format, and accuracy rewards based on IoU and L1 distance, to stimulate robust reasoning in segmentation contexts.
Although Seg-Zero demonstrates strong performance in emergent reasoning tasks, its complete decoupling of the reasoning model and segmentation decoder prevents access to pixel-level feedback from the segmentation results, thereby increasing the risk of reward hacking. To address this, involving the segmentation decoder directly in the reward loop, as a reward provider, not only ensures alignment between optimization objectives and task goals but also alleviates the need for extensive human-annotated reasoning data, enabling a more efficient and scalable learning paradigm.
@@
Moreover, we introduce a subtle modification to the clipped objective of PPO to fully utilize its potential in the reasoning segmentation task. First, we increase the upper clipping threshold to encourage updates from highly advantageous actions, thereby granting the model greater flexibility in optimizing the task-specific reasoning model. Second, we observe that GRPO may occasionally produce overly lengthy responses with limited informative content. During the GRPO optimization process, overly long responses will confuse the model and prevent it from obtaining higher reward signals, which can lead to reward hacking. Rather than constraining each token in a single response, we treat all tokens within a response group, encouraging the policy model to focus on generating responses with higher information density. By integrating task-specific, fine-grained rewards with a tailored optimization objective, SAM-R1 precisely interprets complex instructions and accurately localizes segmentation targets. Using only 3K training samples, our method surpasses the base model by 34.1% on the challenging ReasonSeg benchmark in zero-shot setting.
In conclusion, our contributions can be summarized as below:
@@ 2.1 MLLMs for Vision and Reasoning Segmentation
Multimodal Large Language Models (MLLMs) have significantly advanced visual understanding, extending from foundational tasks like image captioning and visual question answering [1, 6] to more intricate, fine-grained perception challenges such as image segmentation. A notable direction is reasoning segmentation [39, 2, 27], which necessitates that models interpret implicit user queries and perform logical deduction to generate pixel-level masks. Seminal works like LISA [14] demonstrated the viability of MLLMs for such tasks by interfacing them with segmentation models via specialized tokens. However, these initial approaches frequently depended on Supervised Fine-Tuning (SFT) using datasets with simple categorical labels or rudimentary descriptions [21]. This reliance often curtailed out-of-domain generalization and lacked explicit, interpretable reasoning processes [21, 30], thereby motivating the exploration of methods to instill more robust reasoning capabilities within MLLMs for segmentation.
@@ 3.2.1 DeepSeek R1-Zero and GRPO
$\biggl{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left(\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\right)A_{i}\right)-\beta\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{ref}\right)\right)\biggr{]},$
@@ 3.2.2 Our Training Algorithm
$\biggl{[}\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigl{(}\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\right.$
@@ 4.1 Experimental Setup
We use Qwen2.5VL-7B [1] as our base model and SAM2-Large [25] as the segmentation model. All experiments are conducted on 8$\times$×A100 GPUs. During training, we sample 8 responses per question, set $\varepsilon_{\text{high}}=0.3$, and use a learning rate of $1.0\times 10^{-6}$. To ensure the model’s robustness across different domains, we resize all images to 840$\times$840 before feeding them into the MLLM during both training and evaluation. We follow previous works [14, 21] and use both cIoU and gIoU as evaluation metrics. gIoU is defined by the average of all per-image intersection-over-unions, while cIoU is defined by the cumulative intersection over the cumulative union.
@@ 4.3 Main Results
ReasonSeg. Table 1 shows the zero-shot performance of SAM-R1 on the ReasonSeg benchmark. Our method achieves 60.2% gIoU and 54.3% cIoU on the test set, outperforming the previous best, Seg-Zero (58.3% gIoU and 53.4% cIoU). This improvement is mainly due to our fine-grained reward design, which integrates SAM into the RL loop to provide IoU-based feedback during training, aligning reasoning with segmentation. Unlike Seg-Zero’s decoupled design, our unified framework introduces finer-grained segmentation rewards, enabling stable optimization and better generalization with only 3k training samples. Additionally, our improved GRPO strategy—with asymmetric clipping and token-level loss normalization—enhances informativeness and robustness under domain shifts, supporting SAM-R1’s strong zero-shot performance in complex reasoning segmentation. Seg-Zero* denotes performance based on provided model weights, as their reported results used different weights per metric and could not be reproduced.
@@
Referring Expression Segmentation.
Our evaluation results on the Referring Expression Segmentation datasets are shown in Table 2. We use the testA subsets of RefCOCO and RefCOCO+ as OOD test sets, and the test set of RefCOCOg as the in-domain test set. It can be seen that our model, trained on only 3,000 samples, still achieves competitive performance compared to prior methods. Specifically, on the in-domain dataset RefCOCOg, our algorithm SAM-R1 is only 0.2 points lower than Seg-Zero, despite using fewer style-consistent training samples. On the OOD datasets, our model performs comparably to Seg-Zero on RefCOCO, and improves the performance on RefCOCO+ from 73.9 to 74.7. This demonstrates the effectiveness of our approach SAM-R1. We attribute this improvement to the fine-grained reward mechanisms and the flexible exploration strategy, which allows our model to surpass previous out-of-domain performance with significantly less training data.
@@
4.4 Generalization to REC task
Although our model is not trained on any Referring Expression Comprehension (REC) datasets, we observe strong performance on REC task, thanks to the model’s enhanced reasoning ability and fine-grained perceptual capabilities.
As shown in Table 3, our method, SAM-R1, achieves state-of-the-art performance on the LISA-Grounding benchmark with 63.8, significantly surpassing previous methods such as GroundedSAM (26.2), OV-Seg (28.4), X-Decoder (28.5), and Visual-RFT (43.9). This substantial improvement demonstrates the effectiveness of our reinforcement learning-based reasoning framework in complex visual grounding tasks. Unlike prior approaches, which often rely on large-scale supervised training or handcrafted prompt engineering, our method leverages task-aligned rewards and structured reasoning supervision to enable fine-grained object understanding and robust generalization in reasoning-intensive scenarios.
These results demonstrate the generality and adaptability of our method beyond segmentation, highlighting its strong alignment capabilities and transferability to challenging REC scenarios.
Model LISA-Grounding
GroundedSAM 26.2
OV-Seg 28.4
X-Decoder 28.5
Visual-RFT 43.9
SAM-R1(Ours) 63.8
@@ 4.6 Ablation Study
[…]
@@
Ablation Failure: Removing the KL Constraint
@@
Ablation Failure: Encouraging Negative Reference Points
@@
5 Broader Impact
SAM-R1 contributes to the broader goal of making multimodal large models more capable and efficient for complex vision-language tasks. By showing that high-level reasoning can be acquired from weak supervision and task-aligned rewards, our findings highlight the potential for applying similar strategies to other domains, such as video understanding, robotics perception, and medical image analysis, where annotated reasoning data is scarce.
We also observe several limitations that open avenues for future improvement. Although we leverage SAM to provide fine-grained reward signals, its parameters remain frozen throughout training. Future work may explore jointly optimizing SAM alongside the multimodal model to enable deeper alignment between reasoning and segmentation. Additionally, we find that the model struggles to generate meaningful negative reference points, and our current reinforcement learning framework has limited effectiveness in enhancing this ability. Addressing this limitation may further enrich the model’s discriminative reasoning and improve robustness in complex visual scenarios.
6 Conclusion

### v2

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [17, 20, 37, 22, 5, 44, 51] have achieved remarkable progress in the field of visual understanding [36, 42, 27, 43, 12], with their capabilities extending to more complex and fine-grained perception tasks like multimodal segmentation [38, 21, 40]. Compared to conventional segmentation methods that rely on simple categorical labels, the reasoning segmentation task [2, 16, 31] has garnered significant attention for its flexibility and practical applicability, but it also introduces substantially greater challenges. Specifically, it requires models not only to comprehend the intent behind user-provided textual queries accurately but also to perform strong logical reasoning to generate high-quality, pixel-level segmentation outputs.
@@
LISA [16] was the first to introduce the integration of MLLMs with segmentation models via specialized tokens, demonstrating the feasibility of applying MLLMs to reasoning segmentation tasks. Building on this foundation, subsequent studies [25, 2, 31, 41] have adopted similar strategies, leveraging task-specific tokens generated by MLLMs to improve pixel-level segmentation performance. While these approaches are promising, they often rely heavily on large-scale annotated datasets to jointly fine-tune the language model and the segmentation decoder. This not only increases training costs but also raises the risk of catastrophic forgetting, where models perform well on in-domain data but fail to generalize to out-of-domain scenarios [7]. Furthermore, the reasoning segmentation tasks frequently involve ambiguous and complex text queries from users, which demand strong reasoning capabilities from MLLMs to accurately interpret intent and precisely localize the target segmentation regions.
Recent research has shown that reinforcement learning (RL) can significantly enhance the reasoning capabilities of large language models (LLMs) through reward-based feedback mechanisms [13]. DeepSeek-R1 [9] leverages rule-based rewards to further improve the model’s capacity for complex reasoning. This method requires the model to undergo an extensive reasoning process before producing a final answer, with rewards assigned solely based on the correctness of the final response and its adherence to a predefined output format. Such rule-based reward designs align naturally with visual understanding tasks, which often come with accurate ground-truth (GT) annotations. Inspired by this, numerous efforts [8, 46, 50, 39] have applied Group Relative Policy Optimization (GRPO) [33] to vision-language models, incorporating task-specific reward signals. For example, VLM-R1 [34] introduces both format and accuracy rewards for general vision-language tasks, and further incorporates customized rewards tailored to specific applications to mitigate reward hacking. Seg-Zero [23] expands this paradigm by designing a more comprehensive reward system, including reasoning-format, segmentation-format, and accuracy rewards based on IoU and L1 distance, to stimulate robust reasoning in segmentation contexts. Although Seg-Zero demonstrates strong performance in emergent reasoning tasks, its complete decoupling of the reasoning model and segmentation decoder prevents access to pixel-level feedback from the segmentation results, thereby increasing the risk of reward hacking. To address this, involving the segmentation decoder directly in the reward loop, as a reward provider, not only ensures alignment between optimization objectives and task goals but also alleviates the need for extensive human-annotated reasoning data, enabling a more efficient and scalable learning paradigm.
@@
Moreover, we introduce a subtle modification to the clipped objective of PPO to fully utilize its potential in the reasoning segmentation task. First, we increase the upper clipping threshold to encourage updates from highly advantageous actions, thereby granting the model greater flexibility in optimizing the task-specific reasoning model. Second, we observe that GRPO may occasionally produce overly lengthy responses with limited informative content. During the GRPO optimization process, overly long responses will confuse the model and prevent it from obtaining higher reward signals, which can lead to reward hacking. Rather than constraining each token in a single response, we treat all tokens within a response group, encouraging the policy model to focus on generating responses with higher information density. By integrating task-specific, fine-grained rewards with a tailored optimization objective, SAM-R1 precisely interprets complex instructions and accurately localizes segmentation targets. Using only 3K training samples, our method surpasses the base model by 34.1% on the challenging ReasonSeg benchmark in zero-shot setting. In conclusion, our contributions can be summarized as below:
@@ 2.1 MLLMs for Vision and Reasoning Segmentation
Multimodal Large Language Models (MLLMs) have significantly advanced visual understanding, extending from foundational tasks like image captioning and visual question answering [1, 26, 6] to more intricate, fine-grained perception challenges such as image segmentation. A notable direction is reasoning segmentation [45, 2, 31], which necessitates that models interpret implicit user queries and perform logical deduction to generate pixel-level masks. A relevant line of research [10, 35, 11] focuses on using a single generic prompt to perform segmentation, thereby reducing the reliance on manually provided, image-specific inputs. Seminal works like LISA [16] demonstrated the viability of MLLMs for such tasks by interfacing them with segmentation models via specialized tokens. However, these initial approaches frequently depended on Supervised Fine-Tuning (SFT) using datasets with simple categorical labels or rudimentary descriptions [23]. This reliance often curtailed out-of-domain generalization and lacked explicit, interpretable reasoning processes [23, 34], thereby motivating the exploration of methods to instill more robust reasoning capabilities within MLLMs for segmentation.
@@ 3.2.1 DeepSeek R1-Zero and GRPO
$\hskip-50.00008pt\biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left(\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\right)A_{i}\right)-\beta\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{ref}\right)\right)\biggr],$
@@ 3.2.2 Our Training Algorithm
$\hskip-10.00002pt\biggl[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigl(\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\right$.
@@ 4.1 Experimental Setup
[…]
@@ 4.3 Main Results
ReasonSeg. Table 1 shows the zero-shot performance of SAM-R1 on the ReasonSeg benchmark. Our method achieves 60.2% gIoU and 54.3% cIoU on the test set, outperforming the previous best, Seg-Zero (58.3% gIoU and 53.4% cIoU). This improvement is mainly due to our fine-grained reward design, which integrates SAM into the RL loop to provide IoU-based feedback during training, aligning reasoning with segmentation. Unlike Seg-Zero’s decoupled design, our unified framework introduces finer-grained segmentation rewards, enabling stable optimization and better generalization with only 3k training samples. Additionally, our improved GRPO strategy—with asymmetric clipping and token-level loss normalization—enhances informativeness and robustness under domain shifts, supporting SAM-R1’s strong zero-shot performance in complex reasoning segmentation. Seg-Zero-7B* denotes performance based on provided model weights, as their reported results used different weights per metric and could not be reproduced.
@@
Referring Expression Segmentation. Our evaluation results on the Referring Expression Segmentation datasets are shown in Table 2. We use the testA subsets of RefCOCO and RefCOCO+ as OOD test sets, and the test set of RefCOCOg as the in-domain test set. It can be seen that our model, trained on only 3,000 samples, still achieves competitive performance compared to prior methods. Specifically, on the in-domain dataset RefCOCOg, our algorithm SAM-R1 is only 0.2 points lower than Seg-Zero, despite using fewer style-consistent training samples. On the OOD datasets, our model performs comparably to Seg-Zero on RefCOCO, and improves the performance on RefCOCO+ from 73.9 to 74.7. This demonstrates the effectiveness of our approach SAM-R1. We attribute this improvement to the fine-grained reward mechanisms and the flexible exploration strategy, which allows our model to surpass previous out-of-domain performance with significantly less training data.
@@ 4.5 Ablation Study
GRPO
57.8 51.2 74.1 71.8
Token-level
58.0 51.7 74.5 72.4
Clip higher
59.1 52.8 74.9 72.5
Ours
60.2 54.3 75.4 73.1
4.6 Generalization to REC task
Model LISA-Grounding
GroundedSAM 26.2
OV-Seg 28.4
X-Decoder 28.5
Visual-RFT 43.9
SAM-R1(Ours) 63.8
Although our model is not trained on any Referring Expression Comprehension (REC) datasets, we observe strong performance on REC task, thanks to the model’s enhanced reasoning ability and fine-grained perceptual capabilities. As shown in Table 5, our method, SAM-R1, achieves state-of-the-art performance on the LISA-Grounding benchmark with 63.8, significantly surpassing previous methods such as GroundedSAM (26.2), OV-Seg (28.4), X-Decoder (28.5), and Visual-RFT (43.9). This substantial improvement demonstrates the effectiveness of our reinforcement learning-based reasoning framework in complex visual grounding tasks. Unlike prior approaches, which often rely on large-scale supervised training or handcrafted prompt engineering, our method leverages task-aligned rewards and structured reasoning supervision to enable fine-grained object understanding and robust generalization in reasoning-intensive scenarios. These results demonstrate the generality and adaptability of our method beyond segmentation, highlighting its strong alignment capabilities and transferability to challenging REC scenarios.
4.7 Broader Impact and Discussion
Our work shows that reinforcement learning, guided by a segmentation model, can effectively cultivate reasoning in multimodal models. The strong performance of SAM-R1 with only 3,000 training samples highlights a promising path toward data efficiency. By using standard segmentation masks as the supervisory signal, our approach bypasses the need for costly and potentially biased, manually annotated reasoning chains, thus enhancing scalability. More broadly, this study supports a paradigm where models learn complex reasoning from task-aligned rewards rather than explicit instructions. This shift toward learning from weaker, accessible supervision is particularly impactful for domains with scarce reasoning data, such as robotic perception and medical image analysis.
We recognize several limitations for future work. First, SAM’s parameters remain frozen, creating a one-way information flow that prevents it from adapting to the reasoning model. Jointly optimizing both models is a compelling next step. Though computationally demanding, this could foster a synergistic alignment where the models co-adapt. Second, our model struggles to generate meaningful negative reference points, a key capability for robust discriminative reasoning. Our RL framework failed to encourage this, suggesting a foundational limitation that may require new architectural or algorithmic solutions to improve robustness in complex visual scenes.
5 Conclusion
6 Acknowledgements
This work was supported by the STI 2030-Major Projects under Grant 2021ZD0201404.
Appendix A Technical Appendices
@@
A.1 Ablation Failure: Removing the KL Constraint
@@
A.2 Ablation Failure: Encouraging Negative Reference Points
@@
A.3 Failure Analysis
As illustrated in Figure 5, we present several typical failure cases, which mainly highlight two issues: incomplete segmentation and over-segmentation.
A notable observation is that our SAM-R1, through its Thinking process (fourth column), successfully comprehends the prompt and correctly localizes the target object(s). However, this correct semantic understanding does not always translate perfectly into the final segmentation mask. For example, in cases of incomplete segmentation: In the first row, the model correctly identifies the fence, but the Predict mask only covers a small portion of the target. In the second row, the model recognizes The fire extinguishers but incorrectly segments only one of the two instances. Conversely, over-segmentation is demonstrated in the third row: The prompt asks for a part of the car (the hood), and the Thinking process also pinpoints the open hood. However, the model incorrectly segments the entire vehicle instead of just the specified part.
These examples indicate that, while our model performs well in high-level semantic reasoning, limitations still exist in its ability to precisely map this understanding to pixel-level masks, particularly concerning fine-grained segmentation and instance completeness. This remains a key area for future improvement.
Figure 5: Visualization of some failure cases for our SAM-R1 method on the ReasonSeg-val dataset, which shows that our approach still has some limitations.
A.4 Data Efficiency and Scalability Analysis
To investigate the scalability and data efficiency of SAM-R1, we conducted additional experiments by increasing the size of the training data from 3k to 10k. The results clearly show that our method is highly data-efficient, with performance saturating at just 3k samples.
We present the direct comparison in Table 6. As shown, increasing the data to 10k results in negligible fluctuations in ReasonSeg: the cIoU shifts slightly from 55.8 to 55.5 on the val split and from 54.3 to 53.9 on the test split. Similarly, on the RefCOCO benchmarks, we observe only marginal gains, which strongly indicates that performance has already plateaued.
Table 6: Data efficiency analysis with 3k vs. 10k training samples.
Method ReasonSeg RefCOCO Benchmarks
Val Test refCOCO refCOCO+ refCOCOg
SAM-R1 (Ours, 3k) 55.8 54.3 79.2 74.7 73.1
SAM-R1 (Ours, 10k) 55.5 53.9 79.9 75.3 73.5
Gain -0.3 -0.4 +0.7 +0.6 +0.4
From these results, it is evident that our method’s core performance saturates at 3k samples. Given the substantial increase in training cost versus the minimal performance returns, we deliberately chose 3k samples as the optimal trade-off point for demonstrating our method’s capabilities.